I have a problem that reads as follows: The total length of two ships is 2,351 feet. The difference in their lengths is 3 feet.
a) Find the length of each ship.
b) If a football field has a length of 100 yards, determine the length of the longer ship in terms of number of football fields.
The length of the longer ship is ____ ft.
The length of the shorter ship is ____ ft.
The length of the longer ship is _____ football fields.

Answer by mananth(16949)

x+y=2351
x-y =3
add both equations
2x=2354
x= 1177 feet
1174 feet is the second ship

number of football fields = 1177/300
3.92 times